how do you figure out a a formula for a sequence than whats the ninth term, Ex 2 4 6 8
1
Expert's answer
2010-11-12T04:33:28-0500
It's the aruthmetic progression, a[n] = a1 +(n-1)*b, where b = 2, a1 = 2;
so
a[9] = 2 + (9-1)*2 = 18.